OpenStudy (anonymous):

A two-digit number from 10 to 99, inclusive, is chosen at random. What is the probability that this number is divisible by 5? Please note: I don't know how to find the numbers divisible by 5 easily without trying it manually! Is there an easier way to solve it?

OpenStudy (inkyvoyd):

Well, you simply divide the range by 5...

OpenStudy (inkyvoyd):

From 10 to 99, there are (99-10)+1 numbers.

OpenStudy (inkyvoyd):

Starting from 10 (which is divisible by 5), you keep going up.

OpenStudy (anonymous):

I don't understand. =|

OpenStudy (inkyvoyd):

From 1 to 95, there are 95/5 numbers divisible by 5. 95/5=19

OpenStudy (anonymous):

well there 10-15-20-25-30-35-40-45-50-55-60-65-70-75-80-85-90-95 so it would be 18/ 88 can be divided by 5; which can be simplified to 9/44

OpenStudy (anonymous):

well theres*

OpenStudy (inkyvoyd):

Anyways, the number of numbers divisible by 5 from 1 to 95 is 19. But we are taking the number from 10 to 99, so we would have to subtract 1 number, to get 18 numbers from 10 to 99 that are divisible by 5.

OpenStudy (anonymous):

*****my mistake it would be 18/89 can be divided by 5***

OpenStudy (anonymous):

I don't understand the subtraction part. Why do we need to subtract one with the range 10-99?

OpenStudy (inkyvoyd):

Because, for 1-99 there are 19 numbers divisible by 5. But, we started with 10, and not 1. The number 5, which is between 1 and 10, has been counted in that 19. So, since we don't include the 5, we have to subtract it to get 18.

OpenStudy (anonymous):

Does that mean I have to find the multiples of five within the range before solving it?

OpenStudy (inkyvoyd):

No. You only have to determine the starting point and finish point, and essentially you could do 1000-9999 inclusive. The work would not be much more tedious if you use my method, although for 10-99 you can just list out all the numbers.

OpenStudy (anonymous):

I prefer your method. But could you please explain it again from the beginning?

OpenStudy (inkyvoyd):

Well, take 99-10=89. Then, divide by 5 to get 17R4. Since we started with a number divisble by 5 (we started with 10), we add 1 to our answer to get 18. Next, find the number of numbers from 10 to 99, including 10 and 99. To do this, take 99-10+1., which is 90. To see why we have to add one number, consider the number of numbers from 1 to 5, with 2 and 5 included. We have 2,3,4,5, which is 4 numbers, or 5-2+1. We have to add the extra 1 after the 5-2. Next, take 18 and divide by 90 to get 18/90. Simplify the fraction.

OpenStudy (inkyvoyd):

I'll give you two more examples with larger numbers, but I won't explain them out throuorlhy

OpenStudy (anonymous):

What's 17R4? =)

OpenStudy (inkyvoyd):

A two-digit number from 17 to 9959, inclusive, is chosen at random. What is the probability that this number is divisible by 5? Number of numbers divisible by 5: (9959-17)/5=1988R2. Because 9959 and 17 aren't divisible by 5, we don't add anything to 1988. Next, we find the total number of numbers from 17 to 9959. Take 9959-17+1, because we are counting the ends. 9959-17+1=9943 Finally, divide both numbers. Answer:1988/9943

OpenStudy (inkyvoyd):

By the way, I meant a number, not a 2 digit nuimber.

OpenStudy (inkyvoyd):

A number from 12 to a one million minus 2, inclusive, is chosen at random. What is the probability that this number is divisible by 3? (999998-12)/3=333328R2. Both 999998 and 12 are divisible by 3 (the sum of their digits is divisible by 3). Thus, we add two to the number to get 333330. (999998-12)+1=999,987. So, we take 333330/999,987 and simplify the fraction.

OpenStudy (inkyvoyd):

Does that help?

OpenStudy (anonymous):

from the first example i'm still blanked out on this part: Next, we find the total number of numbers from 17 to 9959. Take 9959-17+1, because we are counting the ends. 9959-17+1=9943 I don't understand it.

OpenStudy (inkyvoyd):

Okay. "Find the total number of numbers from 5 to 10, inclusive" we have 5 6 7 8 9 10, which is 6 numbers. In other words, we take 10-5+1=6

OpenStudy (anonymous):

Thank you for your time. I finally get it now haha sorry for the trouble!

OpenStudy (inkyvoyd):

No problem - it took me a very long time before I figured out these kinds of problems on my own. Thanks for taking the time to read :)

OpenStudy (anonymous):

Of course =)